Myesha bought stock in a company two years ago that was worth xx dollars. During the first year that she owned the stock, it decreased by 29%. During the second year the value of the stock increased by 36%. Write an expression in terms of xx that represents the value of the stock after the two years have passed.
Answer by ikleyn(52904)

.
Myesha bought stock in a company two years ago that was worth xx dollars. During the first year that
she owned the stock, it decreased by 29%. During the second year the value of the stock increased by 36%.
Write an expression in terms of xx that represents the value of the stock after the two years have passed.
~~~~~~~~~~~~~~~~~~


The value of the stock after 2 years is  (1+0.36)*(1-0.29)*xx = 1.36*0.71*xx = 0.9656*xx  dollars.    ANSWER

Solved.